OpenStudy (anonymous):

i have to find the perimeter of a square but the only the hypotenuse is given...its 48m....can anyone help?

OpenStudy (anonymous):

 \[x^2+x^2=48^2\] \[2x^2=2304\] \[x^2=1152\] And here is your answer :)